40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 10 02:53:32 2014, Jenne, Update, LSC, Locking activities - nothing new :( LocklossZoom.png
    Reply  Wed Sep 10 02:59:42 2014, ericq, Update, LSC, Locking activities - nothing new :( 
Message ID: 10486     Entry time: Wed Sep 10 02:59:42 2014     In reply to: 10485
Author: ericq 
Type: Update 
Category: LSC 
Subject: Locking activities - nothing new :( 

Some small things I did tonight which did little to nothing to help:

  • I reset the offsets in the SQRTINV FMs to try and match the DC level of the ALS CARM error signal as best as possible, to avoid moving away from the set-point too much, as I was worried we were wandering into regions of too low optical gain. 
  • I turned off the WFS, and hand tweaked the MC alignment. The WFS loops / matrices definitely have some room for improvement, and I was worried that excess angular motion of the MC was coupling into CARM. MC refl is much calmer in the last ~1.5 hrs since I turned off the WFS. 

My main concern with tonights situation was the huge low frequency fluctuations of TRY while CARM/DARM locked on ALS. We saw this being very smooth very recently, but when one arm is fluctuating by multiple line widths, it isn't surprising that locks aren't stable. I want to know why the out of loop stability is so unpredictable. 

ELOG V3.1.3-